Posts

Showing posts from November, 2025

Projection Close 11/22

To do this evening/ was completed: More big crowds for a big, big day of Wicked 2. Made the rounds and checked in with Chenelle about the event tomorrow and on Tuesday. Since the BluRay of Guardians is still at WB, I reached out to Kat to see if she could rip it Monday morning and put it on the Share drive for us to DCP.  Finally got that DCP of Home Alone ingesting, which flew by once it finished all the digest stuff. Took a hot minute, but it'll be available for tomorrow. Please take a look at this on screen first thing in the morning. Thank you! I am ONLY ingesting this to T7, not the LMS.  Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Looks like the lens needs to be shifted upwards slightly as I noticed Bugonia bleeding a little on the bottom ma...

Projection Close 11/21

To do this evening/ was completed: Big crowds for Wicked - still sounding good at 5.5  Inspected all of Dark Angel , which is back on shipping reels tails out and ready for house reels when the time comes. Really great colors. Overall a very nice print.  Predators, for the private event, wasn't fully copied over to the share drive, so I copied over the rest of the files and was able to successfully ingest it. Additionally, today I received the keys for this feature and ingested them. There's a set for Monday and a set for Tuesday (first week of December).  Home Alone DCP still running in T1. Should be fine for Sunday.  Saw there was a package for Alec at the MOD station. Brought it up to 4/6.  Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleani...

NPP Projection Report 11/22 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ripped the Death Race 2000 blu-ray for DCP conversion on the T1 Mac. Once Home Alone finishes, which'll be tonight, we can get this one started. No rush, but I'd like to save this Home Alone DCP as it's a likely pick for future private events. Please leave it on George Ivan and I'll upload it to Google Drive tomorrow. Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: I reached out to DMS for an updated Screenwriter license as the current one will be expiring next Sunday. Gregg is on it. Swept and mopped in 4/6 Cue sheets updated/checked (every shift): I added tomorrow's private event to the cue sheet for tomorrow. It had Frankenstein in the first round for T7. Marquee updated?: Posters updated?: Cleaning l...

NPP Projection Admin 11/21

To do this evening/ was completed: Playlists built for Cuckoo's Nest, Wild Goose Lake, Nightmare on Elm Street (still needs slide). Wild Goose Lake ingested to LMS. Looked at Legends of the Fall, all good. Tech'd all of the NoBudge48 shorts with Desmond, all good, volume and aspect ratio details are listed in the event memo. Put Nightmare on house reels. Ripped the Home Alone blu-ray and started converting the file in T1, it should be done tomorrow evening. If we think the 4/6 desktop will be significantly faster then we can start another conversion later this afternoon. T1 booth swept and mopped. Charlie Chocolate Factory slide ingested. Bugonia print picked up, Dark Angel print dropped off, resting in T1. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning:...

NPP Projection Close 11/20

To do this evening/ was completed: Presentation Issues: Kept an ear on it this evening and  Wicked 2 sounds fine at 5.5 even with the larger crowds Preshow Clips/Playlist: Built all first run CCAP and OCAP show playlists for the week of 11/28 and saved them to the LMS + updated  Zootopia 2 SPLs and removed its trailer/replaced it with the  Spongebob trailer Built and scheduled a pre-show for Spartan on 11/25 (downloaded and converted the intro slide as well)  Built a show playlist for East of Eden on 11/29-30 which is in 2.55:1 so this will have some very minor letterboxing in the 2.39 lens macro...additionally there are no credits so I have put a minute of black at the end of the feature to keep the lights up before the cleaning lights are triggered Trailer Updates: Special Events/ AV set up: B'twixt Now and Sunrise screened this evening in theater 4 without issue Ordering/Receiving (posters & trailers): Teaser posters for The Super Mario Galaxy Movie , The Devi...

NPP Projection Report 11/20 Admin

To do this evening/ was completed: Wicked sounds better at 5.5, so I'm keeping it there for now. Please feel free to monitor this tonight for bigger crowds! - L Presentation Issues: Preshow Clips/Playlist: Just a note on new A1s - please be sure to not put the trailer for the same title in its playlist, even if it comes out a day or two earlier. I updated Eternity with the correct trailer information so it did not have its own trailer playing, a long with Hamnet , and added in the Romance preshow per Desmond and Kat's note. - L Built the playlists for Hamnet and scheduled both this and Eternity on the LMS for Wednesday and Thursday. - L  Trailer Updates: Ingested Marty Supreme v2, Precode Parade_Dec25, Avatar F&A v3, Is This Thing On v2 rev, Spongebob v4 - L Special Events/ AV set up: Masking updates in T3 and T7. I did some tinkering with the cues, both 3 and 7 now have Scope, 2.00, 1.85, 1.78, and 1.66 masking cues plus a "Variable Ratio" cue which is current...

NPP Projection Close 11/19

To do this evening/ was completed: Film Feast went smoothly on my end, mic broken down (they opted not to use it). Finished downloading the NoBudge files and have converted three so far to the share drive: GEZA, When I Get Home, This Is Nuts. Ingested Hamnet to LMS. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 11/19 Admin

To do this evening/ was completed: Staff screenings are out! Blocked on the calendar, scheduled on the LMS. - L Cleaned the platter and ran the first and only round of Bugonia on 35mm today. Our last one! Breaking it down tomorrow with Alec. - L Presentation Issues: There was a dowser glitch at the start of Jay Kelly in T4, first round. The dowser open cue fired but the dowser did not open, as occasionally happens. All it took was hitting the dowser close and open button and rewinding the movie back to the beginning. It was about 3 minutes in when this was caught. A/L Preshow Clips/Playlist: Updated the Goodfellas with the new 5min preshow ingested today. - L Trailer Updates: Special Events/ AV set up: Checked Goodfellas for tonight and set up a mic + stand in T3. All set for the film feast tonight. -A Ordering/Receiving (posters & trailers):  Ordering/Receiving (content): Ingested East of Eden . - L I noticed that, when looking ahead at the LMS schedule, the Bugonia show ...

NPP Projection Close 11/18

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Legends of the Fall intro and screening went very well (I did not have a moment to double check its pre-show as I was busy setting up for trivia so I missed seeing the Dark Angel trailer was absent from it...I played this manually after Cristina and her guest made their remarks) Cleaned both projectors following the show and ran the RP-40 loop for Bugonia  tomorrow morning/refocused the lens/taped the port glass/set the top roller back in place...closed both exhaust ports and put the mic is back in the booth in its charging port Had a hell of a time getting trivia ready since people were there as early as 2PM in order to get seats and by the time I started to set the table and wiring and everything else up the room was nearly full. Today would have benefited from an even earlier set up if I am being completely honest but I digress. Alec had set up XLRs beneath...

NPP Projection Admin 11/18

To do this evening/ was completed: Looked at Barry Lyndon and Btwixt, volumes and lights up put in playlists. 1.66 lens file slightly keystoned in T2, Alec said he'd tweak it tomorrow. **Did not look at Goodfellas, so check on that tonight or tomorrow morning.** -I sorted the 1.66 lens file for T2. It just needed a tiny bit of cropping at the bottom corners and to be zoomed in, it's now perfectly flush with the masking -A Zootopia and Eternity playlists built and scheduled. Ran loop on on P2 for tonight's show. P1 projector cleaned before this afternoon's Bugonia screening, which went... 11/28 forecasts/trailer list hasn't come in yet so not able to schedule 11/28 - 12/4 yet. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: For Trivia tonight, Blank Check podcast is joining and they want to plug into our mixer to record audio. I've got this set up, there are two XLR cables under the AV bench in Trees. They are bringing a Z...

NPP Projection Close 11/17

To do this evening/ was completed: T3 and T7 events went smoothly. T7 pushed back the intro a few times so there was almost some direct overlap but it worked out fine. Wrapped up the mics and cable after each show concluded. Scheduled the coming week in LMS, some first run Wed/Thurs stuff still needs to be built, will be soon. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Admin 11/17

To do this evening/ was completed: Note for Alec/Jack - if Jack is up for it, the lamp could get changed in T7 tomorrow.  Will ran first round of Bugonia  on 35mm. Projector cleaned. Presentation Issues:  Preshow Clips/Playlist: The Crow playlist tech'd. All set for tonight - Lil Trailer Updates: Special Events/ AV set up: Went through the playlists for the special events tonight and both are ready to go. Mics/stands are ready in T3. Spots remote in the basket.  T7 - playlist checked and I brought a mixer down from T6 with some xlr cables and mics. Set everything up. Will brought down chairs. I found that the spots remote didn't seem to be working, so I changed the battery but that didn't help, so I got the ladder out and changed the batteries in the actual spotlights and that did the trick. - Lil Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Received Eternity - ingesting to LMS. Ingested Zootopia to LMS. - Lil Note: the kdms for Barry...

NPP Projection Close 11/16

To do this evening/ was completed: Bugonia went smoothly, projector cleaned afterward. New brunch letters up. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Open 11/16

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Desmond had to cancel this morning's intro to Kiki's Delivery Service  and asked that the slide be removed so I rescheduled the pre-show from yesterday morning. Ran loops prior to the first round. Everything went smoothly and the print is tails out in booth 1. Desmond mentioned holding on to the print so I am leaving it there for the time being. Cleaned both projectors and refocused projector 1/moved the top roller back into place for this afternoon's screening of Bugonia  Turned up the feature volume for  The Doors during the first round. Otherwise the film played without any issues. I removed the feature content from its server/the LMS following the show Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets...

Projection Close 11/15

To do this evening/ was completed: Had a great last (for me, at least!) show of Bugonia  with a huge crowd. Lovely to see so many people tonight.  Presentation Issues: No issues.  Preshow Clips/Playlist: Double checked the playlists for 11/14 (SE) this week and checked the new 11/21 playlists both SE and A1.  Built the playlist for the private event on 11/17. Scheduled on the LMS.  Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Keys will open up for The Crow tomorrow.  Outgoing/Shipping (content): Got rid of some content we don't need on certain servers.  Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Platter and projector have been cleaned.  Cue sheets updated/checked (every shift): Yes.  Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 11/15 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Kiki's Delivery Service went well in T1. I'd recommend playing it at 7.5 on both processors, full house today and tomorrow. I added a volume cue for the CP750 to the playlist. Both projectors have been cleaned and the print is rewound onto house reels for tomorrow. Ran a loop to do the tape masking for Bugonia before the second round in T1. A heads up for Tuesday: Blank Check podcast is doing trivia and might be recording the audio. If so, I'll make sure it's set up during the day and they'll only have to plug in and hit record. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by e...

NPP Projection Close 11/14

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Second round screening of Bugonia  in 35mm went smoothly...cleaned the projector/platter/rollers etc thoroughly and moved the top roller for tomorrow morning's 35mm show (port glass tape can be adjusted once loops are run) Ordering/Receiving (posters & trailers): Additional Wicked: For Good  posters were delivered and stored behind 3/7  Ordering/Receiving (content): B'Twixt Now and Sunrise , Barry Lyndon , and GoodFellas  keys were delivered this evening via Deluxe Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Double checked all cues/schedule for tomorrow (note: theaters 5 and 7 have very short flip times so anticipate some calls for house lights while the staff cleans) Marquee updated?: Posters updated?: Added a "now showing" snipe ab...

Projection Admin 11/14

To do this evening/ was completed: We had Jack's review today. Had a good talk about all things projection related. Alec and I have good feedback for what we can work on department wise, too.    Between the Mountain and the Sky tech all good, volume and lights up logged (5.0, 1:30:00). All A1 11/21 playlists built, plus two of the three SE's for the week. No Other Choice posters in 3/7. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Private event for Frankenstein in T3 went well today. Hosts showed a little late, but everything was perfectly fine. - L Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested Hellboy 2, Crow keys came in! (they open up this Sunday, so I will set a reminder, but most likely I'll take a look on Monday), DCP'd and ingested the Scott Barley Intro for Dark Room - L Outgoing/Shipping (content): Print/video Inspection: Went over the Dark Room shorts playlist and intro video wit...

NPP Projection Close 11/13

Image
To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: John Woods's video game launch party event went down sans  Godspeed You Black Emperor (this was swapped for Akira on the Plex at Trees)...the playlist link was not saved to the regular Spotify account so there was a bit of troubleshooting to sign in and out and back in again to the projection one in order to have the proper music playing for this...everything went smoothly otherwise Bugonia 35mm screening in theater 1 this evening went well (the print continues to look and sound as good as ever) and the machines/platters/rollers etc were cleaned thoroughly following the show Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Beneath the Mountain and the Sky DCP was transferred to the LMS and theater 7 this evening Hellboy II: The Golden Army  DCP was delivered via Deluxe this evening Outgoing/Shipping (content): Print/video Inspect...

Projection Admin 11/13

Image
To do this evening/ was completed: No issues with T3 today. Checked in this morning and the first round has gone just fine. See notes below in maintenance. - L Presentation Issues: Preshow Clips/Playlist: Lots of playlist updates. Got rid of old content/playlists we don't need. Built all SE playlists for 11/14 Kiki's, The Doors, The Crow, Barry Lydon, Btwixt Now ad Sunrise, Dark Room Shorts (which, yes, have a q&A after each short!), Legends of the Fall and Goodfellas including the A1 Wicked 2 playlists, but we're waiting on that Xmas Book trailer. I reached out to Kat. Update: Xmas Book trailer has been ingested. Playlists updated. - L Built just about  Trailer Updates: Special Events/ AV set up: Built and scheduled the playlist for the school screening tomorrow morning. All set - L Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested Wicked 2, Wicked 15min_V2 preshow, transferred The Doors feature to T3, Ingested The Crow (now schedul...

NPP Projection Close 11/12

To do this evening/ was completed: Inspected the first five reels of Legends of the Fall. Good condition so far. Legends DCP ingested to LMS. T3 intro and Q&A went smoothly. Dr Moreau intro and print went smoothly. Had to run around a bit to set up mics for both theaters, T4 wasn't set and T3 needed a few more mics connected and ready to go for the Q&A. Equipment broken down afterward. Moreau is rewound back on its shipping reels. Downloaded T3 log file 11/12 - 11/13. It's the most recent file added to 4/6 desktop (log_11-13-2025). Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes   Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 11/12 Lillian

To do this evening/ was completed: Rebooted the digital showtimes display as yesterday's info was still up.  Had our cinema meeting. All good on content, just reached out to Chenelle about the private event on 11/17.  Presentation Issues: We had an issue in T3 where MOD radioed me to say that cleaning lights were on and the film had stopped. We had lost connection to the server. I went down to reboot the projector, resat everything and then booted it up again. Got the server back up and went straight into One Battle as the issue happened just a few minutes into the feature. It's stable right now, but I'm going to reboot the server in between rounds and pull logs. Spoke to Rebecca and Chenelle about making any updates to the private event. This was a quick turnaround, so were only 7 minutes behind schedule for preshow for the private event.  Lil update: private event is on the way now. Jack will pull logs once all shows are done tonight.  Preshow Clips/Playlist: Staff...

NPP Projection Close 11/11

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Made a DCP slide for the  GoodFellas  event and transferred it to the LMS Built staff show playlists for Jay Kelly , Sentimental Value , and  The Running Man ( Jay Kelly timecode information was in the KDM email + added house lights up cues to all of its SPLs for this week and next) Trailer Updates: Special Events/ AV set up: Night Nurse screening went very smoothly. Framing and changeovers were clean throughout. Turned the volume up to 9.0 since it sounded quiet at 7.5. Mics were returned to the booth to charge. Projectors were cleaned and exhaust ports closed. Ran the RP-40 loop between rounds and refocused the 1.85 lens for Bugonia .  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Put Night Nurse  back on shipping reels and scheduled a FedEx pickup for tomorrow (tracking ID: 885962012718) Asset t...

NPP Projection Open 11/11

To do this evening/ was completed: Barry Lyndon, Goodfellas  ingested to LMS. Looked at Dragon and Good Light, volumes/lights added to playlists. Bugonia print show went smoothly. Ran loops for Night Nurse afterward, projectors cleaned. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Cue sheets for next week are set apart from Wicked For Good, The Crow, Barry Lyndon, and a private event in the middle round on Monday for Theater 7 that hasn't been fully confirmed yet.  Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 11/10

To do this evening/ was completed: Both Bugonia shows went smoothly. Projector cleaned. Scheduled the coming week in LMS. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 11/10

To do this evening/ was completed: The Bugonia OCAP for 11/07, which was first round in T2 today, did not execute Bugonia masking because it was not in the playlist. This was the only playlist without it, so it must have been a simple mistake. Please double check your work when building playlists and we will hold ourselves accountable to double check them once they're built. Thank you, Will, for catching this today. - L Die, My Love is all fine to go. - L Presentation Issues: Preshow Clips/Playlist: Built the new A1 playlists for 11/07 and 11/14. Built SE playlist for Come See Me in the Good Light and How To Train Your Dragon (all scheduled now)- L Trailer Updates: Special Events/ AV set up: Checked out Darkman today. All set. - L Kat added the documentary Godspeed, You Black Emperor to the Share drive, which I then added to Plex in a separate folder so it is easy to find. John asked that we have this playing in Trees during the launch party on Thursday. - L Ordering/Receiving...

NPP Projection Close 11/9

To do this evening/ was completed: Both Bugonia shows went smoothly, projectors cleaned per usual. No masking issues with Die, My Love this evening. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Open 11/9

To do this evening/ was completed: Presentation Issues: MOD radioed about a masking issue in theater 7 during  Die My Love  in the second round of shows. This was not as urgent as I thought it would be as the image was slightly pillarboxed from the 1.37 masking cue that does not go fully flush against the 1.33 frame. Will double check this tomorrow but the playlists are fine and I am not entirely sure the curtains go that flat in that room anyway Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Jaws screening in theater 1 this afternoon. Cleaned both machines and refocused projector 1 after the show for  Bugonia later this evening. The port glass is taped to fit the RP-40 loop again. I left the scope lens from Jaws on projector 2 for the time being. Both lenses will need to be swapped for Night Nurse  on 11/11 Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Jaws is...

NPP Projection Report 11/8 Open

To do this evening/ was completed: Thanks to whoever vacuumed under the T1 screen yesterday, it's much appreciated! Presentation Issues: Preshow Clips/Playlist: Jaws had the Precode Parade 5min preshow clip in it for some reason, instead of the Jaws 5min clip. Fixed it before the theater opened. Trailer Updates: Special Events/ AV set up: Jaws went very well. Cleaned both projectors and rewound the print onto house reels. 1.85 lens back in P1 and focused for Bugonia later. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Our new bookshelf came, I'll put it together on my next admin here. Cue sheets updated/checked (every shift): Marquee updated?: Added Jaws to the brunch side. Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 11/7

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Bugonia screenings in theater 1 went smoothly. The print continues to look and sound great. Everything was cleaned thoroughly and powered down following the final round Ordering/Receiving (posters & trailers): Ordering/Receiving (content): The Crow DCP was delivered via Deluxe Outgoing/Shipping (content): Print/video Inspection: Spent most of the evening inspecting A Nightmare on Elm Street 4: The Dream Master in booth 1 and filled out its inspection report/uploaded stills to the shared drive (this print opens with trailers for Lucky Stiff and Tougher than Leather which might be fun to play for the crowd) Asset tracking doc updated/checked?: Maintenance/cleaning:  Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Yes Additional Notes: A closed captioning de...

NPP Projection Admin 11/7

To do this evening/ was completed: Night Nurse on house reels, first reel tech'd, volume logged. Inspected Spartan, excellent condition, on house reels, stills on Drive, emailed Conrad. Presentation Issues: Preshow Clips/Playlist: Happy that I took a look at the preshow content and first runs just to double check them because Die, My Love was extremely loud at 5.5 in quite a few scenes. I adjusted the volume to 4.8, which seems good in both large and small screening rooms. Also, please be extra careful when building playlists with the trailer volumes as the Val trailer/Book_V3 and Legends trailers did not have the proper cues in the playlists for our new first runs (Die My Love, Predator, It Was Just An Accident). I have adjusted all of them. FYI - the volumes are still up on the white board, so please reference it when building. Thanks! - L Tech'd playlist for Real Genius. - L Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Rec...

NPP Projection Mid+Close (Lillian & Will)

To do this evening/ was completed: Presentation Issues: Chenelle left one of the spotlights on in theater 6 following the Below the Clouds private event...I got a call from Wyatt about customers noticing a "yellow light" on the corner of the screen and figured it could only be the spots...went into the auditorium and turned it off by standing directly in front of the screen ( Megalopolis style, much to the chagrin of everyone seated, I'm sure) and the film continued normally Preshow Clips/Playlist: Updated all playlists that needed the Generic Scifi 15min_v2 as we received it from Kat today. (Darkman, Real Genius, Dr. Moreau). Scheduled loads of SE shows on the LMS. - L Trailer Updates: Special Events/ AV set up: Setup the mics in T6 and double checked the event information for tonight: Update - - > the event went well! Started a little late, but there was plenty of time for the Q&A. Cleaned up after! - Lillian Spoons went smoothly and all the gear was put back be...

NPP Projection Report 11/6 Admin

To do this evening/ was completed: During admin tomorrow, someone should vacuum under the T1 screen. We need to be doing this regularly after we have Spoons, as we do tonight. Presentation Issues: Preshow Clips/Playlist: Every Predator Badlands playlist had the 1.37 lens/masking change instead of scope. I corrected them all. Please watch this when building playlists, much easier to just get it right the first time rather than having to update multiple playlists individually after noticing a mistake. We'll need a re-upload of the 15min Wicked preshow clip, Screenwriter says it's corrupt and won't ingest. Kat is away for the weekend and we're not showing this for another two weeks, so we can bug her about it on Monday. Trailer Updates: Special Events/ AV set up: Spoons is set up. Please remember to put both XLR cables for the PA speakers on the hooks beneath the screen when breaking down. There is a hook for the left speaker cable toward that side of the screen. No cab...

NPP Projection Close 11/5

To do this evening/ was completed: Both Bugonia shows went smoothly. Projector surfaces cleaned before/after. Inspected Kiki's Delivery Service, good condition, on house reels. Continued converting shorts: Strawberry Shortcake still going in 4/6. **I started converting this short before I read about the slides for Below the Clouds, which is to say they have not been converted yet, as the DCP is still going at the end of the night. So the playlist will need to be made tomorrow. 6.8 to 7.0 volume for the feature and 1:52:30 lights mid cue during scroll.** The Doors feature ingested to LMS. Kiss Kiss intro went smoothly, mic back in the booth. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters...

Projection Admin 11/05

To do this evening/ was completed: Staff screenings have been built, scheduled and blocked for Thursday and Friday, staying out of any rooms we're teching in. Also, I will be here tomorrow for a mid for event support. - L We are losing Blue Moon , Shelby Oaks , and After the Hunt on Friday. Please make sure the marquee and posters are updated before doors open. -A Presentation Issues: Preshow Clips/Playlist: Built all A1 11/4 playlists (with the exception of Sentimental Value, Running Man, and Jay Kelly, which we do not have yet.) Built the SE playlists for 11/07 with Night Nurse, Jaws, Dr. Moreau, Darkman, Real Genius. NOTE: reached out to Kat about the V2 15min Scifi preshow and she's trying to get it to us as soon as she can, but she's asked that we put the V12 in there for now if she can't get it to us in time, which is what I've done. - L Trailer Updates: Ingested Dark Angel, Home Alone, Charlie and the Chocolate Factory, Hamnet, History of Sound, and the new ...

NPP Projection Close 11/4

Image
To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Brought a mic stand down from theater 1 to theater 3 as requested in the event memo for the Shock-a-Go-Go shorts intro (the spotlight remote is back behind 3/7 and the stand is beneath the curtain in theater 1) Set up for trivia at Trees Lounge...broke everything down following the 35mm show in theater 4 The Color Purple screening + intro in theater 4 went smoothly save for a bad changeover between reels 4 and 5 (reel 4 was mislabeled on the inspection form as having a set of scribe cues when both were grease...I missed the motor cue by a moment longer than was needed to have everything in sync and the crowd saw some leader from the subsequent reel during the changeover ). A similar error occurred between reels 5 and 6 from the splicing of the leader on the latter reel throwing the framing off but this was also quickly fixed The first round 35mm screening of Bugon...

NPP Projection Admin 11/4

To do this evening/ was completed: Ran loops in T4, frame/focus set, projectors cleaned. Mics set in T4 and T3. Looked at Gits, Kiss Kiss  and Below the Clouds, volumes logged.  Gits is 1:37 aspect, so I've updated the playlist. Went through the Below the Clouds DCP with the filmmaker rather precisely. He's decided on volume 6.8 for the feature, 7.0 if it's a full house, with the house lights mid cue firing a little while into the credits (1:52:30). Die, My Love and Predator playlists built, scheduled those in LMS. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 11/3

To do this evening/ was completed: Both Bugonia shows went smoothly. Projector apparatus cleaned before and after shows. It Was Just An Accident and Die, My Love ingested to LMS. Upcoming week mostly scheduled in LMS, some playlists still need to be made.   Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 11/03

To do this evening/ was completed: FYI Bugonia masking in T2 is fine. I transferred the playlists to T2 again from the LMS and that's done the trick. Saving/syncing. - LH Cleared content from all screening rooms and the LMS. - L Presentation Issues: Preshow Clips/Playlist: Built the playlist for The Gits , which still needs to get tech'd. Built the playlist(s) for Predator Badlands . Added Shock A Go Go and Kiss Kiss Bang Bang to the LMS schedule, including Predator . - LH Trailer Updates: Special Events/ AV set up: Blade (1998) is all set for this evening. LH Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Noticed that our first set of keys for Beyond the Clouds (private on 11/06) would not open until 11.05, despite our tech tomorrow. Chenelle asked for an earlier set and they sent them straight away! All good to take a look at it tomorrow. - L Outgoing/Shipping (content): The print of Blade of Fury got picked up today and is on its way to the Musi...

NPP Projection Close 11/2

To do this evening/ was completed: Both Bugonia print shows went smoothly tonight. Cleaned the projector, decks and rollers after each show. Transferred The Gits DCP to LMS, Kiss Kiss Bang Bang to T2, Below The Clouds to T6. Checked cue sheets at end of the night and saw that we still need to make a playlist for Blade tomorrow. Feature itself is already on T2 server. Presentation Issues: Stepped in to T2 last round to see that the masking cue for Bugonia had not fired, so the film was presented pillarboxed. This was towards the end of the film so I didn't attempt to move the curtains. Didn't want to draw attention to it. When I looked at the playback tab in the IMS I didn't see "Bugonia Masking" attached to the black cue that precedes the feature, but that cue is on the macro controls page. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Prin...

NPP Projection Open 11/1

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Added the  Prince of Egypt show playlist to the schedule as it was missing from the first round in theater 3 (removed the feature content from the server following the show) Trailer Updates: Special Events/ AV set up: Ran loops before the  Blade of Fury screening in theater 1 (everything went smoothly) and refocused projector 1 with the RP-40 loop/taped the port glass for Bugonia  between rounds so everything is set for the 35mm shows this evening Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Put  Blade of Fury back on its shipping reels, repackaged it, and scheduled a pickup for tomorrow for it to be circuited to Music Box Chicago (tracking ID: 885683332280) Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Brunch side can be updated ...

Projection Close 11/1

To do this evening/ was completed: Changed out the lamp in T1 yesterday, That sensor warning is fine - it's been like that.  Both rounds of Bugonia went very well tonight. One thing - I noticed that, when taking out the brain to put it onto the next deck, dust and dirt/residue can pop out of the hole it secures itself into, so be sure to check the print out when you do that, take an air can and graze the top of it with air. Also, try to hold a micro fibre towel under the brain when you lift it up. Presentation Issues: Preshow Clips/Playlist: I reached out to Kat for preshow info for Blade .  Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Echoing Alec's note in the previous blog about Blade of Fury needing to get circuited to The Music Box in Chicago. We can send this out first thing Monday. Info on the cans. Print/video Inspection: Asset tracking doc updated/checked?: Mainten...

NPP Projection Report 11/1 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist:  Trailer Updates: Special Events/ AV set up: Nashville had a great show. The print is packed up and ready to go out on the cart, I scheduled a pickup for Monday (a Ground return label came with this print, so Innocent Blood-- and Blade of Fury if that goes out on Monday -- will need to be a separate Express pickup). Projectors have been cleaned. Ran our loop on P1 to focus the 1.85 lens and tape the window for Bugonia after the brunch show. Should be all set for tonight's 35mm shows. Remember to run loops for 1.85 tomorrow before Sundays on Fire. Projector #1 will be focused from  Bugonia shows but #2 won't because Nashville is scope. Also you'll have to re-do the tape masking for both projectors. Also a good idea to run our RP-40 loop on P1 to do the tape masking for Bugonia after SoF, there should be enough time. Ordering/Receiving (posters & trailers): Ordering/Receiving (content)...